OpenStudy (anonymous):

The tangent at a point P on the curve x^4 y^5 = a^9 meet the x and y axes in A and B respectively Then AP:PB=?

OpenStudy (anonymous):

@shubhamsrg dy/dx = -4y/5x

OpenStudy (anonymous):

@shubhamsrg

OpenStudy (shubhamsrg):

let it cut the axes at (a,0) and (0,b) then b/a = 4y/5x  Let the ratio be k:1 then x = a/(k+1) and y = bk/(k+1) => x/y = a/bk => k= ay/bx we know ay/bx =5/4 hence k = 5/4